OpenStudy (anonymous):

After sailing 15 miles, a sailor changed direction and increased the boat's speed by 2mph. An additional 19 miles was sailed at the increased speed. The total sailing time was 4 hours. FIND THE RATE OF THE BOAT FOR THE FIRST 15 MILES. Is this formula correct that I set up? 15/r=19/2r+2=4 hours I know that the time being the same I can set them equal to one another

OpenStudy (anonymous):

sorry didn't mean to repeat the question

OpenStudy (anonymous):

I have my formula corrected in this second one

OpenStudy (dumbcow):

ok so you have the correct equation you just need to solve for r combine fractions on left side cross multiply add like terms i think you will get a quadtratic so use quadratic formula

OpenStudy (anonymous):

Ok I got 15(2r)=19r+4r^2=8r^2?? Is that right?

OpenStudy (dumbcow):

hmm not quite it should be r+2 not 2r 15/r + 19/r+2 = 4

OpenStudy (anonymous):

sorry I thought because he increased by 2 miles per hour that it was going to be times how can I tell the difference?

OpenStudy (dumbcow):

if i said i doubled my speed then i went from 10mph to 20mph i multiplied by 2 if i said i increased my speed by 4mph, then i went from 10mph to 14 mph i added 4

OpenStudy (anonymous):

ohhh

OpenStudy (anonymous):

thanks dumbcow

OpenStudy (dumbcow):

your welcome after you combine fractions you should get 15(x+2) + 19x = 4x(x+2) oh i used x instead of r

OpenStudy (anonymous):

Yes, thank you for making sure I got the quadractic right. I stumble sometimes when I am combining

OpenStudy (anonymous):

d = rt 15 = xt let x be the first rate 19 = (x + 2)t solve both for t 15 -- = t x 19 ----- = t x + 2 We know their combined time was 4 hours.. so we will add these together to get 4 15 19 -- + ----- = 4 x x+2 common denominator of x(x + 2) you multiply every term by that x(x+2)15 x(x+2)19 x(x+2)4 -------- + -------- = x x + 2 cancel the bottoms into the tops (x + 2)15 + x(19) = 4x(x+2) 15x + 30 + 19x = 4x^2 + 8x 34x + 30 = 4x^2 + 8x 0 = 4x^2 - 26x - 30 0 = 2(2x^2 - 13x - 15) 0 = 2(2x - 15)(x + 1) x = 15/2 or -1 since hours can't be negative we use the 15/2 15/2 = 7.5 x = 7.5 mph

OpenStudy (anonymous):

thank you blextiing

OpenStudy (anonymous):

I missed my quadractic like I knew I would putting a 32 where there was supposed be a 36 i mean 26